# Minibatch Optimal Transport and Perplexity Bound Estimation in Discrete Flow Matching

### v3

@@ A.2 First Upper Bound Derivation Details
$\log{p_{t}(x_{1};\theta)}\leq$
$\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-w_{t}^{i}(x^{i},x_{t})\bigg)dt,$ (100)
@@
$\log{p_{t}(x_{1};\theta)}\leq$
$\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\bigg(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\bigg)dt.$ (101)
@@
$H(p_{1},p_{1}(\theta))\leq$
$\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{t},x_{1}}p_{t,1}(x_{t},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\bigg)dt.$ (102)
@@
$H(p_{1},p_{1}(\theta))\leq\mathcal{B}=\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}\bigg($
$\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\bigg)dt,$ (103)
@@ A.3 Alternative Upper Bound Derivation Details
$H(\bar{q}_{1},\bar{p}_{1})\leq H(\bar{q}_{0})-\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})+D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})$
$\sum_{x_{t}}\bar{q}_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(\frac{\bar{q}^{i}_{t}(x^{i}|x_{0},x_{1})}{\bar{q}^{i}_{t}(x_{t}^{i}|x_{0},x_{1})}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{\tilde{w}_{t}^{i}(x_{t}^{i},x)}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})\bigg)dt$ (104)
@@
$H(p_{1},p_{1}(\theta))\leq H(p_{0})-\int_{0}^{1}\sum_{x_{t}}p_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})dt+\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})$
$\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(\frac{p_{t}^{i}(x^{i}|x_{0},x_{1})}{p_{t}^{i}(x^{i}_{t}|x_{0},x_{1})}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{\tilde{w}_{t}^{i}(x_{t}^{i},x)}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)\bigg)dt.$ (105)
@@ A.3.1 Masked source special case
$H(p_{1},p_{1}(\theta))\leq$
$\mathcal{B}=\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\frac{\dot{k}_{t}}{k_{t}}(1-\delta_{m}(x_{t}^{i}))-\delta_{m}(x_{t}^{i})\frac{\dot{k}_{t}}{1-k_{t}}\log{\frac{k_{t}}{1-k_{t}}}$
@@
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k}_{t}}{1-k_{t}}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{x_{t}^{i}\neq x^{i}}+$
$(1-p^{i}_{1|t}(x_{t}^{i},x_{t};\theta))-\delta_{x_{t}^{i}\neq x^{i}}\log{p^{i}_{1|t}(x_{1}^{i},x_{t};\theta)}dt\bigg).$ (123)
@@ A.4 MD4 Special Case
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k}_{t}}{1-k_{t}}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{m}(x_{t}^{i})+$
$\delta_{m}(x_{t}^{i})(1-p^{i}_{1|t}(x_{t}^{i},x_{t};\theta))-\delta_{m}(x_{t}^{i})\log{p^{i}_{1|t}(x_{1}^{i},x_{t};\theta)}dt\bigg),$ (124)
@@
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k}_{t}}{1-k_{t}}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p^{i}_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{m}(x_{t}^{i})p^{i}_{1|t}(x_{t}^{i},x_{t};\theta)+$
$\delta_{m}(x_{t}^{i})\log{p^{i}_{1|t}(x_{1}^{i},x_{t};\theta)}dt\bigg).$ (125)
@@
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k}_{t}}{1-k_{t}}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{m}(x_{t}^{i})\log{p^{i}_{1|t}(x_{1}^{i},x_{t};\theta)}dt\bigg).$ (126)
@@ A.5 The Precise Perplexity
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})=D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+$
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\bigg)dt$-
@@
$\log{p_{t}(x_{1};\theta)}=$
$\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-w_{t}^{i}(x^{i},x_{t})\bigg)dt$-
$\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(r_{p_{t|1}}w_{t}^{i}(x_{t}^{i},x)\log{\frac{r_{p_{t|1}}w_{t}^{i}(x_{t}^{i},x)}{r_{p_{t}^{\theta}}v_{t}^{i}(x_{t}^{i},x)}}+r_{p_{t}^{\theta}}v_{t}^{i}(x_{t}^{i},x)-r_{p_{t|1}}w_{t}^{i}(x_{t}^{i},x)\bigg)dt,$ (128)
@@
$\log{p_{t}(x_{1};\theta)}=$
$\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-w_{t}^{i}(x^{i},x_{t})\bigg)dt$
$H(p_{0})+\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})dt-$ (129)
$\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(\frac{p_{t|1}(x|x_{1})}{p_{t|1}(x_{t}|x_{1})}w_{t}^{i}(x_{t}^{i},x)\log{\frac{w_{t}^{i}(x_{t}^{i},x)}{r_{p_{t}^{\theta}}u_{t}^{i}(x_{t}^{i},x;\theta)}}+r_{p_{t}^{\theta}}u_{t}^{i}(x_{t}^{i},x;\theta)\right)dt.$ (130)
Taking the expectation with respect to the data distribution
$H(p_{1},p_{1}(\theta))=-H(p_{0})+$
$\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg[\tilde{w}_{t}^{i}(x^{i},x_{t})+$
$\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-w_{t}^{i}(x^{i},x_{t})\bigg)$ (131)
$\left(\frac{p^{i}_{t|1}(x^{i}|x_{1},x_{0})}{p^{i}_{t|1}(x^{i}_{t}|x_{1},x_{0})}w_{t}^{i}(x_{t}^{i},x)\log{\frac{w_{t}^{i}(x_{t}^{i},x)}{r_{p_{t}^{\theta}}u_{t}^{i}(x^{i},x_{t};\theta)}}+r_{p_{t}^{\theta}}u_{t}^{i}(x^{i},x_{t};\theta)\right)\bigg]dt.$ (132)
The only terms above that we do not have an explicit form of are the learned-probability ratios between neighbor states. These are the terms missing if we tried to directly calculate the loglikelihood at a point using the continuity equation,
@@ A.6 Time-independence of predictive probabilities in Multimasked flows
$p^{i}_{t|0,1}(z^{i}|x_{0},x_{1})=(1-k_{t})\delta_{x^{i}_{0}}(z^{i})+k_{t}\delta_{x^{i}_{1}}(z^{i})=\begin{cases}(1-k_{t})\delta_{x^{i}_{0}}(z^{i}),&z^{i}\leq V,\\ k_{t}\delta_{x^{i}_{1}}(z^{i}),&z^{i}>V.\end{cases}$ (134)
@@
$p_{t|0,1}(z\mid x_{0},x_{1})=\Biggl[\prod_{i:\,z^{i}>V}(1-k_{t})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V}k_{t}\Biggr]\Biggl[\prod_{i:\,z^{i}>V}\delta_{x^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V}\delta_{x^{i}_{1}}(z^{i})\Biggr]$ (135)
@@
$\frac{\Bigg(\Biggl[\prod_{i:\,z^{i}>V}(1-k_{t})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V}k_{t}\Biggr]\Bigg)\Biggl[\prod_{i:\,z^{i}>V}\delta_{x^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V}\delta_{x^{i}_{1}}(z^{i})\Biggr]p(x_{0},x_{1})}{\sum_{\tilde{x}_{0},\tilde{x}_{1}}\Bigg(\Biggl[\prod_{i:\,z^{i}>V}(1-k_{t})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V}k_{t}\Biggr]\Bigg)\Biggl[\prod_{i:\,z^{i}>V}\delta_{\tilde{x}^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V}\delta_{\tilde{x}^{i}_{1}}(z^{i})\Biggr]p(\tilde{x}_{0},\tilde{x}_{1})}$ (137)
$\frac{\Biggl[\prod_{i:\,z^{i}>V}\delta_{x^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V}\delta_{x^{i}_{1}}(z^{i})\Biggr]p(x_{0},x_{1})}{\sum_{\tilde{x}_{0},\tilde{x}_{1}}\Biggl[\prod_{i:\,z^{i}>V}\delta_{\tilde{x}^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V}\delta_{\tilde{x}^{i}_{1}}(z^{i})\Biggr]p(\tilde{x}_{0},\tilde{x}_{1})}=p_{0,1}(x_{0},x_{1}\mid z)$ (138)

### v4

@@ A.2 First Upper Bound Derivation Details
$\log{p_{t}(x_{1};\theta)}\leq\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-w_{t}^{i}(x^{i},x_{t})\bigg)dt,$ (97)
@@
$\log{p_{t}(x_{1};\theta)}\leq\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\bigg(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\bigg)dt.$ (98)
@@
$H(p_{1},p_{1}(\theta))\leq\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{t},x_{1}}p_{t,1}(x_{t},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\bigg)dt.$ (99)
@@
$H(p_{1},p_{1}(\theta))\leq\mathcal{B}$ (100)
where
$\mathcal{B}:=\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}\bigg(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\bigg)dt,$ (101)
@@ A.3 Alternative Upper Bound Derivation Details
$H(\bar{q}_{1},\bar{p}_{1})\leq H(\bar{q}_{0})-\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})+D_{KL}(\bar{q}_{0}\|\bar{p}_{0})$
$\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}\bar{q}_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(\frac{\bar{q}^{i}_{t}(x^{i}|x_{0},x_{1})}{\bar{q}^{i}_{t}(x_{t}^{i}|x_{0},x_{1})}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{\tilde{w}_{t}^{i}(x_{t}^{i},x)}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})\bigg)dt$ (102)
@@
$H(p_{1},p_{1}(\theta))\leq H(p_{0})-\int_{0}^{1}\sum_{x_{t}}p_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})dt$
$\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(\frac{p_{t}^{i}(x^{i}|x_{0},x_{1})}{p_{t}^{i}(x^{i}_{t}|x_{0},x_{1})}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{\tilde{w}_{t}^{i}(x_{t}^{i},x)}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)\bigg)dt.$ (103)
@@ A.3.1 Masked Source Special Case
$H(p_{1},p_{1}(\theta))\leq\mathcal{B}=\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\frac{\dot{k}_{t}}{k_{t}}(1-\delta_{m}(x_{t}^{i}))-\delta_{m}(x_{t}^{i})\frac{\dot{k}_{t}}{1-k_{t}}\log{\frac{k_{t}}{1-k_{t}}}$
@@
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k}_{t}}{1-k_{t}}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{x_{t}^{i}\neq x^{i}}+(1-p^{i}_{1|t}(x_{t}^{i},x_{t};\theta))-\delta_{x_{t}^{i}\neq x^{i}}\log{p^{i}_{1|t}(x_{1}^{i},x_{t};\theta)}dt\bigg).$ (121)
@@ A.4 MD4 Special Case
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k}_{t}\ dt}{1-k_{t}}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{m}(x_{t}^{i})+\delta_{m}(x_{t}^{i})(1-p^{i}_{1|t}(x_{t}^{i},x_{t};\theta))-\delta_{m}(x_{t}^{i})\log{p^{i}_{1|t}(x_{1}^{i},x_{t};\theta)}\bigg)$ (122)
@@
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k}_{t}}{1-k_{t}}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p^{i}_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{m}(x_{t}^{i})p^{i}_{1|t}(x_{t}^{i},x_{t};\theta)-\delta_{m}(x_{t}^{i})\log{p^{i}_{1|t}(x_{1}^{i},x_{t};\theta)}\bigg)dt.$ (123)
@@
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k}_{t}}{1-k_{t}}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}p_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\bigg(-\delta_{m}(x_{t}^{i})\log{p^{i}_{1|t}(x_{1}^{i},x_{t};\theta)}\bigg).$ (124)
@@ A.5 The Precise Perplexity
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})=D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\bigg)dt$
@@
$\log{p_{t}(x_{1};\theta)}=\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-w_{t}^{i}(x^{i},x_{t})\bigg)dt-$
$\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(r_{p_{t|1}}w_{t}^{i}(x_{t}^{i},x)\log{\frac{r_{p_{t|1}}w_{t}^{i}(x_{t}^{i},x)}{r_{p_{t}^{\theta}}v_{t}^{i}(x_{t}^{i},x)}}+r_{p_{t}^{\theta}}v_{t}^{i}(x_{t}^{i},x)-r_{p_{t|1}}w_{t}^{i}(x_{t}^{i},x)\bigg)dt,$ (126)
@@
$H(p_{1},p_{1}(\theta))=H(p_{0})+\int_{0}^{1}\mathbb{E}_{p_{0,1}(x_{0},x_{1})}\mathbb{E}_{p_{t|0,1}(x_{t}|x_{0},x_{1})}\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t}|x_{1})dt$
$\int_{0}^{1}\mathbb{E}_{p_{0,1}(x_{0},x_{1})}\mathbb{E}_{p_{t|0,1}(x_{t}|x_{0},x_{1})}\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(u_{t}^{i}(x^{i},x_{t}|x_{1})\log{\frac{u_{t}^{i}(x^{i},x_{t}|x_{1})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-u_{t}^{i}(x^{i},x_{t}|x_{1})\bigg)dt$
$\int_{0}^{1}\mathbb{E}_{p_{0,1}(x_{0},x_{1})}\mathbb{E}_{p_{t|0,1}(x_{t}|x_{0},x_{1})}\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(\frac{p^{i}_{t|1,0}(x^{i}|x_{1},x_{0})}{p^{i}_{t|1,0}(x_{t}^{i}|x_{1},x_{0})}u_{t}^{i}(x_{t}^{i},x|x_{1})\log{\frac{u_{t}^{i}(x_{t}^{i},x|x_{1})}{\frac{p^{\theta}_{t}(x)}{p^{\theta}_{t}(x_{t})}u_{t}^{i}(x_{t}^{i},x;\theta)}}+\frac{p^{\theta}_{t}(x)}{p^{\theta}_{t}(x_{t})}u_{t}^{i}(x_{t}^{i},x;\theta)\bigg)dt.$ (127)
The only terms above that we do not have an explicit form of are the learned-probability ratios between neighbor states. These are the terms missing if we tried to directly calculate the log-probability at a point using the continuity equation,
@@ A.6 Time-independence of Predictive Probabilities in Multimasked Flows
$p^{i}_{t|0,1}(z^{i}|x_{0},x_{1})=(1-k_{t})\delta_{x^{i}_{0}}(z^{i})+k_{t}\delta_{x^{i}_{1}}(z^{i})=\begin{cases}(1-k_{t})\delta_{x^{i}_{0}}(z^{i}),&z^{i}>V_{d},\\ k_{t}\delta_{x^{i}_{1}}(z^{i}),&z^{i}\leq V_{d}.\end{cases}$ (129)
@@
$p_{t|0,1}(z\mid x_{0},x_{1})=\Biggl[\prod_{i:\,z^{i}>V_{d}}(1-k_{t})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V_{d}}k_{t}\Biggr]\Biggl[\prod_{i:\,z^{i}>V_{d}}\delta_{x^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V_{d}}\delta_{x^{i}_{1}}(z^{i})\Biggr]$ (130)
@@
$\frac{\Bigg(\Biggl[\prod_{i:\,z^{i}>V_{d}}(1-k_{t})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V_{d}}k_{t}\Biggr]\Bigg)\Biggl[\prod_{i:\,z^{i}>V_{d}}\delta_{x^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V_{d}}\delta_{x^{i}_{1}}(z^{i})\Biggr]p(x_{0},x_{1})}{\sum_{\tilde{x}_{0},\tilde{x}_{1}}\Bigg(\Biggl[\prod_{i:\,z^{i}>V_{d}}(1-k_{t})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V_{d}}k_{t}\Biggr]\Bigg)\Biggl[\prod_{i:\,z^{i}>V_{d}}\delta_{\tilde{x}^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V_{d}}\delta_{\tilde{x}^{i}_{1}}(z^{i})\Biggr]p(\tilde{x}_{0},\tilde{x}_{1})}$ (132)
$\frac{\Biggl[\prod_{i:\,z^{i}>V_{d}}\delta_{x^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V_{d}}\delta_{x^{i}_{1}}(z^{i})\Biggr]p(x_{0},x_{1})}{\sum_{\tilde{x}_{0},\tilde{x}_{1}}\Biggl[\prod_{i:\,z^{i}>V_{d}}\delta_{\tilde{x}^{i}_{0}}(z^{i})\Biggr]\Biggl[\prod_{i:\,z^{i}\leq V_{d}}\delta_{\tilde{x}^{i}_{1}}(z^{i})\Biggr]p(\tilde{x}_{0},\tilde{x}_{1})}=p_{0,1}(x_{0},x_{1}\mid z)$ (133)